# Axi-majoron for almost everything

### v2

@@ 4.2 Spontaneous leptogenesis via axi-majoron
$Y_{B,\rm sb}^{>}$$\approx$ $\frac{0.54\times c_{B-L}b_{N}\lambda_{\phi}^{5/4}}{\pi^{1/2}x_{\rm fo}^{2}g_{*,\rm d}^{1/4}{\rm B}_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{h}^{2}}\right)^{1/2}Y_{\Phi,e}$ (4.38)
$\simeq$ $1.8\times 10^{-10}\left(\frac{b_{N}}{{\rm B}_{1}}\right)\left(\frac{\xi_{\phi}}{200}\right)^{5/4}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{Y_{\Phi,e}}{0.1}\right)\ ,$
where we have used $x_{\rm fo}=10$ and $c_{B-L}=2$ in the second line. The observed baryon asymmetry can be obtained by interplay between $b_{i},\ \xi_{\phi},\ m_{\nu_{1}}$, and $Y_{\Phi,e}$. Comparing Eq. (4.38) to Eq. (4.18), we see that the contribution via spontaneous leptogenesis mechanism can not catch up the one from AD-mechanism in this case, too.
@@ 4.3 Resonant leptogenesis
If the inflaton does not carry enough amount of asymmetry, the baryon number asymmetry of the present universe should be obtained by decays of RHNs, following the usual way of out-of-equilibrium decay. Because of the tiny $y_{\nu}$ in our scenario, the so-called ’resonant leptogenesis’ [30] may be the only possibility in this case. The essential point of the mechanism is that the CP violation in the decay of $N_{i}$ mass-eigenstate can be given by
@@
and $\left(\Delta m\right)_{ij}\equiv|m_{i}-m_{j}|$. Hence, if $\left(\Delta m\right)_{ij}\sim\Gamma_{i}\simeq\Gamma_{j}$, it is possible to have $\delta_{i}\sim\mathcal{O}(1)$. Since the lightest RHN with $\left|y_{\nu_{1}}\right|\ll\left|y_{\nu_{2,3}}\right|$ is to recover the standard thermal bath via its delayed decay (see Section 3.3), two heavier RHN mass-eigenstates, $N_{2,3}$, are responsible for the resonant leptogenesis. Lepton number asymmetry is produced well before the electroweak phase transition. It is diluted by the later entropy injection from the decay of the long-lived lightest RHN, $N_{1}$. The present baryon number asymmetry is then expected to be [30]
@@
$Y_{B,\rm res}$ $\sim$ $\frac{\pi^{2}{\tilde{\delta}}_{23}c_{m}^{2}}{700\left(c_{m}^{2}+1\right)\lambda_{\phi}^{1/4}{\rm B}_{1}\sqrt{{\rm B}_{N}}}\left(\frac{g_{*}(T_{\gamma,N})}{15\sqrt{2}}\right)^{1/2}\frac{\sqrt{m_{\nu_{1}}\phi_{0}}v_{h}}{m_{\nu}M_{\rm P}}$ (4.43)
$\simeq$ $\frac{3.0\times 10^{-6}\times{\tilde{\delta}}_{23}c_{m}^{2}}{\left(c_{m}^{2}+1\right){\rm B}_{1}\sqrt{{\rm B}_{N}\xi_{\phi}}}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}\ .$
@@
Figure 12: The case of resonant leptogenesis for the late time baryon number asymmetry, $Y_{B,\rm res}$ in Eq. (4.43) matching observations as a function of $m_{\nu_{1}}$ and ${\rm b_{1}}$ for ${\tilde{\delta}}_{23}=2\times 10^{-5}$(solid red), $4\times 10^{-5}$(dashed red) wtih $c_{m}=1,\ b_{N}=0.15$, $\xi_{\phi}=50$, and $\phi_{0}=\phi_{0}^{\rm ref}$. The scheme of the blue region and the blue lines is the same as the one in Fig. 10.
@@ 5 Dark matter and dark radiation
In our scenario, two types of axi-majorons appear: hot and cold axi-majorons. Hot axi-majorons are produced in decays of the Peccei-Quinn inflaton as discussed in Section 3. At the epoch of the late time decay of the lightest RHN mass-eigenstate, the energy density of axi-majorons to the radiation density of SM particles is given by
@@
where the subscript ‘0’ stands for quantities at the present universe, $\rho_{\nu}$ is the energy density of a single neutrino species, $T_{\gamma,\rm d}$ and $T_{\gamma,0}$ are respectively the photon temperatures at the epoch of $N_{i}$’s decay and the present time, and we used $g_{*}(T_{\gamma,\rm d})=g_{*S}(T_{\gamma,\rm d})=106.75$ and $g_{*S}(T_{\gamma,0})=3.9$ in the last line. It may help to ameliorate the tension in observations for the expansion rate of the present universe [32].
@@
Cold axi-majorons in our scenario are produced only via the misalignment mechanism, since $U(1)_{\rm PQ}$ is already broken during inflation and never restored. It is the only component of cold dark matter in our scenario. In the minimal DFSZ axion scenario we are considering here, the axion decay constant is given by $f_{a}=\phi_{0}/c_{a}$ with $c_{a}=6$ being the QCD-anomaly coefficient of $U(1)_{\rm PQ}$ symmetry with the charge of PQ-field $\Phi$ being normalized to be unity. In conventional scenarios where the initial kinetic energy of the axion is ignored, for $\theta_{\rm mis}=\mathcal{O}(1)$ as the misalignment angle, a right amount of dark matter is obtained for $f_{a}\approx 3\times 10^{11}{\rm GeV}$ [75]. If axi-majoron has a sufficiently large kinetic energy to delay the onset of its low energy coherent oscillation due to angular misalignment, there would be a chance of the kinetic misalignment [76] even for $\phi_{0}$ much smaller than $\phi_{0}^{\rm ref}$. It requires $Y_{\Phi}=\mathcal{O}(10)\times\left(\phi_{0}/10^{9}{\rm GeV}\right)$ as the ratio of the PQ-number density to the entropy density of the SM particles around the epoch of axi-majoron oscillations in conventional scenarios. However, in our scenario, as shown in Eq. (4.5), the angular motion is limited as $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ with $\alpha\lesssim\mathcal{O}(10^{-2})$ which is necessary to obtain sufficiently large motion of axi-majoron after inflation (see Eq. (4.5)). Hence, the kinetic misalignment does not work, and the symmetry-breaking scale is constrained as $\phi_{0}\sim\phi_{0}^{\rm ref}$ for $\mathcal{O}(1)$ misalignment angle.
@@
The amount of dark radiation is determined by the duration of the matter-domination (MD) era driven by the lightest RHN, as shown in Eq. (5.2). The presence of an early MD-era reduces the $e$-folds associated with cosmic scales relative to the case of the standard cosmology. As a result, the spectral index in our scenario is expected to be smaller than in the standard case. The difference relative to the expectation in the standard cosmology is small, but this implies that, if some amount of dark radiation is detected in future experiments, in principle, it can be used for a consistency check of our scenario with the spectral index of density perturbations, although it will strongly depend on the sensitivity of the experiments. Also, such an early MD-era causes dilution of pre-existing particles due to the entropy release at the end of the era. Specially, the nearly scale-invariant inflationary gravitational waves would be diluted for cosmic scales which were inside the horizon at the epoch of the decay of the dominating particle. If detected, firstly, the inflationary gravitational waves provide information of the energy scale during inflation, i.e., $H_{I}$ in Eq. (3.20). When combined with inflationary observables, $\lambda_{\phi}$ and $\xi_{\phi}$ can be nearly fixed, modulo the minor variation due to the possible angular motion during inflation when $\alpha$ is somewhat sizable. Additionally, the spectra of the primordial gravitational waves at the present universe will have spectral changes at two frequencies: The characteristic frequencies at the present universe are given by
@@
where $T_{0}$ is the background temperature of the present universe. In the second line of Eq. (5.4), we used $f_{N_{1},\rm d}=H_{\rm d}/2\pi$ with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$, and $H_{\rm ew}$ is the expansion rate of the universe at the epoch of the EWPT with photon temperature $T_{\rm ew}=100{\rm GeV}$. These frequencies are associated with the beginning (Eq. (5.3)) and the end(Eq. (5.4)) of the matter-domination era. All modes of $f>f_{{\rm eq},0}$ are suppressed by the factor $\left(H_{\rm d}/H_{\rm eq}\right)^{2/3}$ while modes of $f<f_{N_{1},0}$ are not affected. Once such a feature of GWs is detected in some future experiments sensitive to the energy scale of inflation in our scenario(corresponding to $\Omega_{\rm GW}h^{2}\sim\mathcal{O}(10^{-17})$ as the fractional energy of GWs at the present universe), it would provide information of $H_{1,\rm eq}$ and $\Gamma_{N_{1}}$ separately [77], as long as $\Gamma_{N_{1}}/H_{1,\rm eq}$ is not much smaller than unity. Even in the case only $f_{N_{1},0}$ is detected, if available, $\Delta N_{\rm eff}$ can be used to determine $H_{\rm eq}$. Then, since dark matter relic density would nearly fix $\phi_{0}$, it may be possible to extract information of $m_{\nu_{1}}$ and $m_{N_{1}}$. To our knowledge, this is a unique way of probing those mass parameters in the seesaw sector.
@@ 6 Conclusions
We showed that our scenario provides a simple unified framework simultaneously addressing the following phenomenological/cosmological/theoretical puzzles: (i) the origin of scales, (ii) a period of primordial inflation, (iii) baryogenesis, (iv) tiny neutrino masses, (v) dark matter, and (vi) strong CP-problem. Primordial inflation in our scenario is realized along the Peccei-Qinn field direction thanks to a non-minimal gravitational interaction and scale-invariance. Thanks to the symmetry-breaking non-minimal coupling, in general angular motion of the inflaton arises, and can generate a large PQ-number asymmetry at the end of inflation. It is nothing but the way of generating particle asymmetry in Affleck-Dine mechanism of baryogenesis, which is naturally realized in our setup. Depending on parameters, the asymmetry can be transferred to the visible sector through right-handed neutrinos in various ways. It can be through a non-thermal decay of the lightest RHNs which have a helicity-asymmetry induced by the asymmetry of PQ-number. This is a kind of usual Affleck-Dine scenario, although details are somewhat different in our scenario. Additionally, the asymmetry in the visible sector can be obtained via thermal processes, as in the case of spontaneous baryogenesis mechanism. In this case, the remnant angular motion of axi-majoron after the decay of the Peccei-Quinn inflaton provides the effective chemical potential to the fields charged under $U(1)_{\rm PQ}$-symmetry. Compared to the contribution by AD-mechanism, this contribution to the baryon number asymmetry is always sub-dominant, though. The third possibility is the resonant leptogenesis from the decay of nearly degenerate heavy RHN mass-eigenstates. This case is relevant when the initial PQ-number asymmetry after inflation is not enough for neither AD-mechanism nor spontaneous baryogenesis.
@@
In regard to axion solution, the symmetry-breaking term in the gravity sector does not invalidate the solution, since its contribution to the axion potential is proportional to the total potential energy of the universe, and becomes negligible in the late universe. Although our scenario is a DSFZ axion model having more than one domain, it does not have a domain wall problem thanks to the sizable non-zero angular motion and weak interactions of Peccei-Quinn field to RHNs. Iso-curvature perturbations of the axi-majoron are also suppressed because of the following reasons: (i) the magnitude of Peccei-Quinn field during inflation is comparable to or larger than the Planck scale, and (ii) the curvature along the angular direction is either positive or much smaller than the expansion rate when the relevant cosmological scales exit the horizon. The axion-quality problem may also be absent in the presence of Gauss-Bonnet term.
Our scenario is a very simple minimal extension of the SM, but it provides a unified picture of the unknown history of the universe from inflation to the BBN epoch, thanks to a scale-invariant $U(1)_{\rm PQ}$-breaking non-minimal gravitational interaction. This model predicts a presence of a short early matter-domination era. It determines the amount of axi-majoron dark radiation, and leaves a characteristic fingerprint on inflationary primordial gravitational waves. If such a feature is detected in future experiments, and once combined with inflationary observables, the amount of dark radiation and the spectra of gravitational waves can provide information of the mass parameters of the lightest states of the left-handed and the right-handed neutrinos. Although complicated, it may end up being the unique way of obtaining those physical quantities.

### v3

@@ 4.2 Spontaneous leptogenesis via axi-Majoron
$Y_{B,\rm sb}^{>}$$\approx$ $\frac{0.54\times c_{B-L}b_{N}\lambda_{\phi}^{5/4}}{\pi^{1/2}x_{\rm fo}^{2}g_{*,\rm d}^{1/4}{\rm B}_{1}}\left(\frac{m_{\nu_{1}}M_{\rm P}}{v_{u}^{2}}\right)^{1/2}Y_{\Phi,e}$ (4.38)
$\simeq$ $\frac{1.8\times 10^{-10}}{\sin\beta}\left(\frac{b_{N}}{{\rm B}_{1}}\right)\left(\frac{\xi_{\phi,0}}{200}\right)^{5/4}\left(\frac{10^{4}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{Y_{\Phi,e}}{0.1}\right)\ ,$
where we have used $x_{\rm fo}=10$ and $c_{B-L}=2$ in the second line. The observed baryon asymmetry can be obtained by interplay between $b_{i},\ \xi_{\phi,0},\ m_{\nu_{1}}$, and $Y_{\Phi,e}$. Comparing Eq. (4.38) to Eq. (4.18), we see that the contribution via spontaneous leptogenesis mechanism cannot catch up the one from AD-mechanism in this case, too.
@@ 4.3 Resonant leptogenesis
If the inflaton does not carry enough asymmetry, the baryon number asymmetry of the present universe should be obtained by decays of RHNs, following the usual way of out-of-equilibrium decay. Because of the tiny $y_{\nu}$ in our scenario, the so-called “resonant leptogenesis” [30] may be the only possibility in this case. The essential point of the mechanism is that the CP violation in the decay of $N_{i}$ mass-eigenstate can be given by
@@
and $\left(\Delta m\right)_{ij}\equiv|m_{i}-m_{j}|$. Hence, if $\left(\Delta m\right)_{ij}\sim\Gamma_{i}\simeq\Gamma_{j}$, it is possible to have $\delta_{i}\sim\mathcal{O}(1)$. Since the lightest RHN with $\left|y_{\nu_{1}}\right|\ll\left|y_{\nu_{2,3}}\right|$ is to recover the standard thermal bath via its delayed decay (see Sec. 3.3), two heavier RHN mass-eigenstates, $N_{2,3}$, are responsible for the resonant leptogenesis. Lepton number asymmetry is produced well before the electroweak phase transition. It is diluted by the later entropy injection from the decay of the long-lived lightest RHN, $N_{1}$. The present baryon number asymmetry is then expected to be [30]
@@
$Y_{B,\rm res}$ $\sim$ $\frac{\pi^{2}{\tilde{\delta}}_{23}c_{m}^{2}}{700\left(c_{m}^{2}+1\right)\lambda_{\phi}^{1/4}{\rm B}_{1}\sqrt{{\rm B}_{N}}}\left(\frac{g_{*}(T_{\gamma,N})}{15\sqrt{2}}\right)^{1/2}\frac{\sqrt{m_{\nu_{1}}\phi_{0}}v_{u}}{m_{\nu}M_{\rm P}}$ (4.43)
$\simeq$ $\frac{3.0\times 10^{-6}\times{\tilde{\delta}}_{23}c_{m}^{2}\sin\beta}{\left(c_{m}^{2}+1\right){\rm B}_{1}\sqrt{{\rm B}_{N}\xi_{\phi,0}}}\left(\frac{10^{3}m_{\nu_{1}}}{m_{\nu}}\right)^{1/2}\left(\frac{\phi_{0}}{\phi_{0}^{\rm ref}}\right)^{1/2}\ .$
@@
Figure 12: The case of resonant leptogenesis for the late time baryon number asymmetry, $Y_{B,\rm res}$ in Eq. (4.43) matching observations as a function of $m_{\nu_{1}}$ and ${\rm b_{1}}$ for ${\tilde{\delta}}_{23}=2\times 10^{-5}$(solid red), $4\times 10^{-5}$(dashed red) wtih $c_{m}=1,\ \sin\beta=1,\ b_{N}=0.15$, $\xi_{\phi,0}=50$, and $\phi_{0}=\phi_{0}^{\rm ref}$. The scheme of the blue region and the blue lines is the same as the one in Fig. 10.
@@ 5 Dark matter and dark radiation
In our scenario, two types of axi-Majorons appear: hot and cold axi-Majorons. Hot axi-Majorons are produced in decays of the Peccei-Quinn inflaton as discussed in Sec. 3. At the epoch of the late time decay of the lightest RHN mass-eigenstate, the energy density of axi-Majorons to the radiation density of SM particles is given by
@@
where the subscript 0 stands for quantities at the present universe, $\rho_{\nu}$ is the energy density of a single neutrino species, $T_{\gamma,\rm d}$ and $T_{\gamma,0}$ are respectively the photon temperatures at the epoch of $N_{i}$’s decay and the present time, and we used $g_{*}(T_{\gamma,\rm d})=g_{*S}(T_{\gamma,\rm d})=106.75$ and $g_{*S}(T_{\gamma,0})=3.9$ in the last line. It may help to ameliorate the tension in observations for the expansion rate of the present universe [32].
@@
Cold axi-Majorons in our scenario are produced only via the misalignment mechanism, since $U(1)_{\rm PQ}$ is already broken during inflation and never restored. It is the only component of cold dark matter in our scenario. In the minimal DFSZ axion scenario we are considering here, the axion decay constant is given by $f_{a}=\phi_{0}/c_{a}$ with $c_{a}=6$ being the QCD-anomaly coefficient of $U(1)_{\rm PQ}$ symmetry with the charge of PQ-field $\Phi$ being normalized to be unity. In conventional scenarios where the initial kinetic energy of the axion is ignored, for $\theta_{\rm mis}=\mathcal{O}(1)$ as the misalignment angle, a right amount of dark matter is obtained for $f_{a}\approx 3\times 10^{11}{\rm GeV}$ [80]. If axi-Majoron has a sufficiently large kinetic energy to delay the onset of its low energy coherent oscillation due to angular misalignment, there would be a chance of the kinetic misalignment [81] even for $\phi_{0}$ much smaller than $\phi_{0}^{\rm ref}$. It requires $Y_{\Phi}=\mathcal{O}(10)\times\left(\phi_{0}/10^{9}{\rm GeV}\right)$ as the ratio of the PQ-number density to the entropy density of the SM particles around the epoch of axi-Majoron oscillations in conventional scenarios. However, in our scenario, as shown in Eq. (4.5), the angular motion is limited as $Y_{\Phi,e}\lesssim\mathcal{O}(1)$ with $\alpha\lesssim\mathcal{O}(10^{-2})$ which is necessary to obtain sufficiently large motion of axi-Majoron after inflation [see Eq. (4.5)]. Hence, the kinetic misalignment does not work, and the symmetry-breaking scale is constrained as $\phi_{0}\sim\phi_{0}^{\rm ref}$ for $\mathcal{O}(1)$ misalignment angle.
@@
The amount of dark radiation is determined by the duration of the matter-domination (MD) era driven by the lightest RHN, as shown in Eq. (5.2). The presence of an early MD-era reduces the $e$-folds associated with cosmic scales relative to the case of the standard cosmology. As a result, the spectral index in our scenario is expected to be smaller than in the standard case. The difference relative to the expectation in the standard cosmology is small, but this implies that, if some amount of dark radiation is detected in future experiments, in principle, it can be used for a consistency check of our scenario with the spectral index of density perturbations, although it will strongly depend on the sensitivity of the experiments. Also, such an early MD-era causes dilution of preexisting particles due to the entropy release at the end of the era. Specially, the nearly scale-invariant inflationary gravitational waves would be diluted for cosmic scales which were inside the horizon at the epoch of the decay of the dominating particle. If detected, firstly, the inflationary gravitational waves provide information of the energy scale during inflation, i.e., $H_{I}$ in Eq. (3.20). When combined with inflationary observables, $\lambda_{\phi}$ and $\xi_{\phi,0}$ can be nearly fixed, modulo the minor variation due to the possible angular motion during inflation when $\alpha$ is somewhat sizable. Additionally, the spectra of the primordial gravitational waves at the present universe will have spectral changes at two frequencies: The characteristic frequencies at the present universe are given by
@@
where $T_{0}$ is the background temperature of the present universe. In the second line of Eq. (5.4), we used $f_{N_{1},\rm d}=H_{\rm d}/2\pi$ with $H_{\rm d}=(2/3)\Gamma_{N_{1}}$, and $H_{\rm ew}$ is the expansion rate of the universe at the epoch of the EWPT with photon temperature $T_{\rm ew}=100{\rm GeV}$. These frequencies are associated with the beginning [Eq. (5.3)] and the end [Eq. (5.4)] of the matter-domination era. All modes of $f>f_{{\rm eq},0}$ are suppressed by the factor $\left(H_{\rm d}/H_{\rm eq}\right)^{2/3}$ while modes of $f<f_{N_{1},0}$ are not affected. Once such a feature of GWs is detected in some future experiments sensitive to the energy scale of inflation in our scenario(corresponding to $\Omega_{\rm GW}h^{2}\sim\mathcal{O}(10^{-17})$ as the fractional energy of GWs at the present universe), it would provide information of $H_{1,\rm eq}$ and $\Gamma_{N_{1}}$ separately [82], as long as $\Gamma_{N_{1}}/H_{1,\rm eq}$ is not much smaller than unity. Even in the case only $f_{N_{1},0}$ is detected, if available, $\Delta N_{\rm eff}$ can be used to determine $H_{\rm eq}$. Then, since dark matter relic density would nearly fix $\phi_{0}$, it may be possible to extract information of $m_{\nu_{1}}$ and $m_{N_{1}}$ if $\tan\beta$ is tightly constrained by collider experiments so as to lead to $\sin\beta\approx 1$ [83]. To our knowledge, this is a unique way of probing those mass parameters in the seesaw sector.
@@ 6 Conclusions
We showed that our scenario provides a simple unified framework simultaneously addressing the following phenomenological/cosmological/theoretical puzzles: (i) the origin of scales, (ii) a period of primordial inflation, (iii) baryogenesis, (iv) tiny neutrino masses, (v) dark matter, and (vi) strong CP-problem. Primordial inflation in our scenario is realized along the Peccei-Qinn field direction thanks to a nonminimal gravitational interaction and scale-invariance. The presence of the symmetry-breaking nonminimal coupling causes in general angular motion of the inflaton, and can generate a large PQ-number asymmetry at the end of inflation. It is nothing but the way of generating particle asymmetry in Affleck-Dine mechanism of baryogenesis, which is naturally realized in our setup. Depending on parameters, the asymmetry can be transferred to the visible sector through right-handed neutrinos in various ways. It can be through a nonthermal decay of the lightest RHNs which have a helicity-asymmetry induced by the asymmetry of PQ-number. This is a kind of usual Affleck-Dine scenario, although details are somewhat different in our scenario. Additionally, the asymmetry in the visible sector can be obtained via thermal processes, as in the case of spontaneous baryogenesis mechanism. In this case, the remnant angular motion of axi-Majoron after the decay of the Peccei-Quinn inflaton provides the effective chemical potential to the fields charged under $U(1)_{\rm PQ}$-symmetry. Compared to the contribution by AD-mechanism, this contribution to the baryon number asymmetry is always subdominant, though. The third possibility is the resonant leptogenesis from the decay of nearly degenerate heavy RHN mass-eigenstates. This case is relevant when the initial PQ-number asymmetry after inflation is not enough for neither AD-mechanism nor spontaneous baryogenesis.
@@
In regard to axion solution, the symmetry-breaking term in the gravity sector does not invalidate the solution, since it is effectively turned off at low energy due to its field-dependent nature which we assumed. The axion-quality problem may also be absent in the presence of Gauss-Bonnet term. Although our scenario is a DSFZ axion model having more than one domain, it does not have a domain wall problem thanks to the sizable nonzero angular motion and weak interactions of Peccei-Quinn field to RHNs that make the symmetry never restored during the whole history of the universe. Isocurvature perturbations of the axi-Majoron are also suppressed because of the following reasons: (i) the magnitude of Peccei-Quinn field during inflation is comparable to or larger than the Planck scale, (ii) the effective curvature along the angular direction is positive and much smaller than the expansion rate when the relevant cosmological scales exit the horizon (i.e., the perturbations of the field orthogonal to the inflaton direction is nearly frozen or suppressed.).
Our scenario is a very simple minimal extension of the SM, but it provides a unified picture of the unknown history of the Universe from inflation to the BBN epoch, thanks to a scale-invariant $U(1)_{\rm PQ}$-breaking nonminimal gravitational interaction. This model predicts a presence of a short early matter-domination era. It determines the amount of axi-Majoron dark radiation, and leaves a characteristic fingerprint on inflationary primordial gravitational waves. If such a feature is detected in future experiments, and once combined with inflationary observables, the amount of dark radiation and the spectra of gravitational waves can provide information of the mass parameters of the lightest states of the left-handed and the right-handed neutrinos. Although complicated, it may end up being the unique way of probing those physical quantities.